The Taste Delight Bakery produces q large chocolate cakes at a cost of $10.24 each and sells (q-3) of them for $20.14 each. The other three cakes are used for display purposes then donated to a charity food bank. The bakery has fixed costs of $2449 per week for rent and $1532 for labour. How many chocolate cakes must it produce per week to break even, assuming the cakes are the only product? (Round your answer to the nearest whole number to the actual break even point.)
Answer #1

The total cost for the bakery per week is

10.24*q+2449+1532

=10.24q+3981

The bakers sells q-3 cakes at a price of 30.14. So, total revenues per week

=(q-3)20.14

To break even, revenue-cost=0. So,

10.24q+3981=(q-3)20.14

q=408.224

Nearest whole number is 408. So,

q=408.